# Dynamic Structural Causal Models

### v1

@@ 1 Introduction
Many real-world systems exhibit a non-trivial dependence on time. In science and engineering, such systems are often modelled with Ordinary Differential Equations when the dynamics are deterministic, with Random Differential Equations (RDEs) when the dynamics of a population is described but every individual can have different (deterministic) dynamics, and with Stochastic Differential Equations (SDEs) when the dynamics of each individual are inherently stochastic (Banks et al., , 2014). In this work, we will focus on the latter formalism, which generalises the former two. SDEs do not come formally equipped with causal semantics. As their mathematical formulation is rather technical, it is a nontrivial task to apply existing methods of causal reasoning to this model class. In this work, we investigate how SDEs can be mapped into a Dynamic Structural Causal Models (DSCMs), originally introduced by Rubenstein et al., (2018) and refined in this work, to naturally bridge this gap. More precisely, we
@@ 2.1 Towards causal reasoning with SDEs
A system of SDEs is assumed to have a certain causal interpretation: for the SDEs in Definition 1, it is assumed that the LHS is determined by variables on the RHS. In Section 4 we will indeed interpret the Itô map $I_{v}$ for and SDE $f_{v}$ as a structural equation of a DSCM. This construction then gives a clear causal interpretation to the graph $G(\mathcal{D})$, where the processes $V$ and $W$ are interpreted as endogenous and exogenous respectively. Compared to existing work where for a systems of constraints one deduces the causal structure (e.g. Iwasaki and Simon, , 1994; Blom et al., , 2021), our setting makes relatively strong assumptions.
@@ 5 Time-evaluations of DSCMs
Figure 3: The graph $G(\mathcal{M})$, time-split $G(\mathcal{M}_{\operatorname*{ev}^{+}(0,s,t)})$ and subsampled $G(\mathcal{M}_{\operatorname*{ev}(0,s,t)})$.
@@
Let a set of distinct time-indices $\tau=\{t_{1},...,t_{m}\}\subseteq\mathcal{T}$ and DSCM $\mathcal{M}=\left\langle V_{p}\times\mathcal{E},W,\mathcal{X}_{V},\mathcal{X}_{W},f,\mathbb{P}(X_{W})\right\rangle$ be given. The time-split DSCM $\mathcal{M}_{\operatorname*{ev}^{+}(\tau)}$ is defined as $\mathcal{M}_{\operatorname*{ev}^{+}(\tau)}=\left\langle\tilde{V},W,\mathcal{X}_{\tilde{V}},\mathcal{X}_{W},\tilde{f},\mathbb{P}(X_{W})\right\rangle$ with as differences from $\mathcal{M}$:
@@
    $\tilde{V}=V_{p}\times\tilde{\mathcal{E}}$, where $\tilde{\mathcal{E}}$ is a further subdivision of $\mathcal{E}$, where for every $t\in\tau$, an interval $[s,u]\in\mathcal{E}$ containing $t$ is partitioned into $\{[s,t),\{t\},(t,u]\}\subseteq\tilde{\mathcal{E}}$; [^11] A similar construction applies for (half-)open intervals.
@@
Given a DSCM $\mathcal{M}$, consider the time-split DSCM $\mathcal{M}_{\operatorname*{ev}^{+}(\tau)}$ with endogenous variables $V\times\mathcal{E}$. Let $\mathcal{E}^{\prime}\subseteq\mathcal{E}$ denote the set of non-singleton intervals. The subsampled DSCM is defined as $\mathcal{M}_{\operatorname*{ev}(\tau)}:=(\mathcal{M}_{\operatorname*{ev}^{+}(\tau)})_{\operatorname*{\mathrm{marg}}(V\times\mathcal{E}^{\prime})}$.
@@
Let $\mathcal{M}_{\mathcal{D}}$ be the DSCM from Example 4, and let $\mathcal{M}:=(\mathcal{M}_{\mathcal{D}})_{\operatorname*{\mathrm{marg}}(L)}$ be the marginalisation of $\mathcal{M}_{\mathcal{D}}$ with respect to $L:=\{X_{3}^{0},X_{3},X_{4}^{0}\}$. The graphs $G(\mathcal{M})$, time-split $G(\mathcal{M}_{\operatorname*{ev}^{+}(s,t)})$ and subsampled $G(\mathcal{M}_{\operatorname*{ev}(s,t)})$ are depicted in Figure 3, where we assume all processes to be temporally Markov. We assume the processes $X_{1},X_{2}$ and $X_{4}$ to be temporally Markov for graphical appeal.
@@
Note that this is consistent with the nomenclature that is earlier introduced, where we referred to a DSCM as collapsed if $\mathcal{E}=\mathcal{T}$.
The following proposition shows that no information is lost when time-splitting a DSCM, but not when subsampling a DSCM.
@@
For a DSCM $\mathcal{M}$ with evaluation index $\mathcal{E}$ that partitions $\mathcal{T}$ and a sequence of increasing time-indices $\tau=(t_{1},...,t_{m})\subseteq\mathcal{T}$, we have $\mathcal{M}=\mathcal{M}_{\operatorname*{ev}^{+}(\tau)*}$. If $\tau\neq\mathcal{T}$, we have $\mathcal{M}\neq\mathcal{M}_{\operatorname*{ev}(\tau)*}$.
@@ 5.1 Local Conditional Independence
In the following theorem, we relate local independence to statistical relation between the variables in time-split DSCMs.
@@ Appendix C Proofs
For a DSCM $\mathcal{M}$ with evaluation index $\mathcal{E}$ that partitions $\mathcal{T}$ and a sequence of increasing time-indices $\tau=(t_{1},...,t_{m})\subseteq\mathcal{T}$, we have $\mathcal{M}=\mathcal{M}_{\operatorname*{ev}^{+}(\tau)*}$. If $\tau\neq\mathcal{T}$, we have $\mathcal{M}\neq\mathcal{M}_{\operatorname*{ev}(\tau)*}$.
@@
The first claim follows directly from Definition 13. For the second claim, let $V\times\mathcal{E}$ and $V\times\mathcal{E}^{\prime}$ denote the endogenous indices of $\mathcal{M}$ and $\mathcal{M}_{\operatorname*{ev}(\tau)*}$ respectively. If $\tau\neq\mathcal{T}$ we have $\mathcal{E}^{\prime}\neq\mathcal{E}$, which proves the result. ∎

### v2

@@ 1 Introduction
Many real-world systems exhibit a non-trivial dependence on time. In science and engineering, such systems are often modelled with Ordinary Differential Equations when the dynamics are deterministic, with Random Differential Equations (RDEs) when the dynamics of a population is described but every individual can have different (deterministic) dynamics, and with Stochastic Differential Equations (SDEs) when the dynamics of each individual are inherently stochastic (Banks et al., , 2014). In this work, we will focus on the latter formalism, which generalises the former two. SDEs do not come formally equipped with causal semantics. As their mathematical formulation is rather technical, it is a nontrivial task to apply existing methods of causal reasoning to this model class. In this work, we investigate how SDEs can be mapped into a Dynamic Structural Causal Model (DSCM), originally introduced by Rubenstein et al., (2018) and refined in this work, to naturally bridge this gap. More precisely, we
@@ 2.1 Towards causal reasoning with SDEs
A system of SDEs is assumed to have a certain causal interpretation: for the SDEs in Definition 1, it is assumed that the left-hand side is determined by variables on the right-hand side. In Section 4 we will indeed interpret the Itô map $I_{v}$ for and SDE $f_{v}$ as a structural equation of a DSCM. This construction then gives a clear causal interpretation to the graph $G(\mathcal{D})$, where the processes $V$ and $W$ are interpreted as endogenous and exogenous respectively. Compared to existing work where for a systems of constraints one deduces the causal structure (e.g. Iwasaki and Simon, , 1994; Blom et al., , 2021), our setting makes relatively strong assumptions.
@@ 5 Time-evaluations of DSCMs
Figure 3: The graph $G(\mathcal{M})$, time-split $G(\mathcal{M}_{\operatorname*{ev}^{+}(0,s,t)})$ and subsampled $G(\mathcal{M}_{\operatorname*{ev}(0,s,t)})$. For graphical appeal, the processes $X_{1},X_{2}$ and $X_{4}$ are assumed to be temporally Markov.
@@
Let a set of distinct time-indices $\tau=\{t_{1},...,t_{m}\}\subseteq\mathcal{T}$ and DSCM $\mathcal{M}=\left\langle V_{p}\times\mathcal{E},W,\mathcal{X}_{V},\mathcal{X}_{W},f,\mathbb{P}(X_{W})\right\rangle$ be given. The time-split DSCM $\mathcal{M}_{\operatorname*{ev}^{+}(\tau)}$ is defined as $\mathcal{M}_{\operatorname*{ev}^{+}(\tau)}=\left\langle\tilde{V},W,\mathcal{X}_{\tilde{V}},\mathcal{X}_{W},\tilde{f},\mathbb{P}(X_{W})\right\rangle$ with differences from $\mathcal{M}$:
@@
    $\tilde{V}=V_{p}\times\tilde{\mathcal{E}}$, where $\tilde{\mathcal{E}}$ is a further subdivision of $\mathcal{E}$, where for every $t\in\tau$, any interval $[s,u]\in\mathcal{E}$ containing $t$ is partitioned into $\{[s,t),\{t\},(t,u]\}\subseteq\tilde{\mathcal{E}}$; [^11] A similar construction applies for (half-)open intervals.
@@
Let $\mathcal{M}_{\operatorname*{ev}^{+}(\tau)}$ be a time-split DSCM with endogenous variables $V_{p}\times\mathcal{E}$. Let $\mathcal{E}^{\prime}\subseteq\mathcal{E}$ denote the set of non-singleton intervals. The subsampled DSCM is defined as $\mathcal{M}_{\operatorname*{ev}(\tau)}:=(\mathcal{M}_{\operatorname*{ev}^{+}(\tau)})_{\operatorname*{\mathrm{marg}}(V_{p}\times\mathcal{E}^{\prime})}$.
@@
Let $\mathcal{M}_{\mathcal{D}}$ be the DSCM from Example 4, and let $\mathcal{M}:=(\mathcal{M}_{\mathcal{D}})_{\operatorname*{\mathrm{marg}}(L)}$ be the marginalisation of $\mathcal{M}_{\mathcal{D}}$ with respect to $L:=\{X_{3}^{0},X_{3},X_{4}^{0}\}$. The graphs $G(\mathcal{M})$, time-split $G(\mathcal{M}_{\operatorname*{ev}^{+}(s,t)})$ and subsampled $G(\mathcal{M}_{\operatorname*{ev}(s,t)})$ are depicted in Figure 3. We assume the processes $X_{1},X_{2}$ and $X_{4}$ to be temporally Markov with time-invariant dynamics for graphical appeal.
@@
Note that this is consistent with the nomenclature that is introduced earlier, where we referred to a DSCM as collapsed if $\mathcal{E}=\mathcal{T}$.
The following proposition shows that no information is lost when time-splitting a DSCM, but information is lost when subsampling a DSCM.
@@
Given a DSCM $\mathcal{M}$ with evaluation index $\mathcal{E}$ that partitions $\mathcal{T}$ and given a set of distinct time-indices $\tau=\{t_{1},...,t_{m}\}\subseteq\mathcal{T}$, we have $\mathcal{M}=\mathcal{M}_{\operatorname*{ev}^{+}(\tau)*}$. On the other hand the map $\mathcal{M}\mapsto\mathcal{M}_{\operatorname*{ev}(\tau)}$ is not invertible, and $\mathcal{M}\neq\mathcal{M}_{\operatorname*{ev}(\tau)*}$ if $\tau\neq\mathcal{T}$.
Analysis of the information-loss in the mapping $\mathcal{M}\mapsto\mathcal{M}_{\operatorname*{ev}(\tau)}$ could be particularly interesting in the light of impossibility results for inference of features of $\mathcal{M}$ from the subsampled DSCM $\mathcal{M}_{\operatorname*{ev}(\tau)}$.
@@ 5.1 Local Conditional Independence
In the following theorem, we relate local independence to statistical relations between the variables in time-split DSCMs.
@@ Appendix C Proofs
Given a DSCM $\mathcal{M}$ with evaluation index $\mathcal{E}$ that partitions $\mathcal{T}$ and given a set of distinct time-indices $\tau=\{t_{1},...,t_{m}\}\subseteq\mathcal{T}$, we have $\mathcal{M}=\mathcal{M}_{\operatorname*{ev}^{+}(\tau)*}$. On the other hand the map $\mathcal{M}\mapsto\mathcal{M}_{\operatorname*{ev}(\tau)}$ is not invertible; in particular $\mathcal{M}\neq\mathcal{M}_{\operatorname*{ev}(\tau)*}$ if $\tau\neq\mathcal{T}$.
@@
The first claim follows directly from Definition 13. For the second claim, let $\mathcal{M}_{1}$ be a DSCM on $\mathcal{T}=[0,1]$ with $V=\{1\}\times\{[0,1]\}$, $W=\emptyset$, and structural equation $X_{1}(t)=\mathbbm{1}_{[0,1/2)}(t)$, and let $\mathcal{M}_{2}$ be the same as $\mathcal{M}_{1}$ but with structural equation $X_{1}(t)=\mathbbm{1}_{[0,1/3)}(t)$. Then for $\tau=\{a,b\}$ with $20\leq a<1/$ 2 and $11/3\leq b<$ 1 we have $(\mathcal{M}_{1})_{\operatorname*{ev}(\tau)}=(\mathcal{M}_{2})_{\operatorname*{ev}(\tau)}$, so the subsampling operation is not invertible. For the final claim let $V\times\mathcal{E}$ and $V\times\mathcal{E}^{\prime}$ denote the endogenous indices of any DSCM $\mathcal{M}$ and $\mathcal{M}_{\operatorname*{ev}(\tau)*}$ respectively. If $\tau\neq\mathcal{T}$ we have $\mathcal{E}^{\prime}\neq\mathcal{E}$, which proves the result. ∎